Problem 4 Find the Present Value at Year 0. (5 points) 1-5% 8 9 10 11 12 13 14 0 1 2 3 4 $3000 5000 $7000 V $9000 $11000
Answer #1

ANSWER:

I = 5%

Present value = cash flow in year 5(p/f,i,n) + cash flow in year 6(p/f,i,n) + cash flow in year 7(p/f,i,n) + cash flow in year 8(p/f,i,n) + cash flow in year 9(p/f,i,n)

Pv = -3,000(p/f,5%,5) - 5,000(p/f,5%,6) - 7,000(p/f,5%,7) - 9,000(p/f,5%,8) - 11,000(p/f,5%,9)

Pv = -3,000 * 0.7835 - 5,000 * 0.7462 - 7,000 * 0.7107 - 9,000 * 0.6768 - 11,000 * 0.6446

Pv = -2,350.5 - 3,731 - 4,974.9 - 6,091.2 - 7,090.6

Pv = -24,238.2

so the present value is -$24,238.2